# Compression and Distillation of Data Quadruplets in Non-intrusive Reduced-order Modeling

### v3

@@ 8 Non-intrusive PORK-based DT-IRKA for Discrete Time Systems
It is now clear that DT-IRKA can be implemented using available transfer function samples $G(\alpha_{i})$ and $G(\beta_{i})$, eliminating the need for repeated measurements of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. The pseudo-code for the PORK-based DT-IRKA (PORK-DTIRKA) is outlined in Algorithm 8.
@@
2:  while(relative change in $\lambda_{i}$ tol)
@@
6:  Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $[b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $[c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
8.1 Tracking the Error $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$
@@
$=$ $||G(z)||_{\mathcal{H}_{2}}^{2}+||G_{r}(z)^{(i-1)}||_{\mathcal{H}_{2}}^{2}-2\text{trace}\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}.$
Thus, with a delay of one iteration, the error $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}$ can be tracked by computing $||G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ in each iteration. Specifically, the variable component of the error in data-driven DT-IRKA can be monitored non-intrusively by tracking the following term:
$||G_{r}(z)^{(i-1)}||_{\mathcal{H}_{2}}^{2}-2\,\text{trace}\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}.$
However, it is important to note that the term $-2\,\text{trace}\Big{(}C_{r}^{(i)}\big{(}C_{r}^{(i-1)}T_{r}^{(i-1)}\big{)}^{*}\Big{)}$ is an approximation and not exact. Its accuracy depends on the precision of the approximation of the integral (73) and (74) or the approximation of the infinite summation (83) and (84).
@@ 9 Compression and Distillation of Data Quadruplets
Throughout this paper, a consistent pattern has emerged in all the discussed data-driven algorithms: each algorithm constructs a Loewner quadruplet (in the frequency domain) or an impulse data quadruplet (in the time domain) and then reduces the respective data quadruplet, as illustrated in Figure 1.
@@
    Similar to numerical integration, interpolatory low-rank BT does not reduce $G(s)$ directly. Instead, an interpolant of $G(s)$ is first implicitly constructed (or explicitly constructed in non-intrusive data-driven settings). This interpolant is not particularly compact, as it is constructed to interpolate $G(s)$ at several interpolation points to capture the majority of the original system’s dynamics. Subsequently, this interpolant acts as a surrogate for $G(s)$. The ROMs produced by these low-rank BT algorithms are approximations of the interpolants of $G(s)$, rather than $G(s)$ itself. In this sense, low-rank BT could be termed “numerical BT” if we wish to adopt terminology analogous to numerical integration.
@@
    QuadBT and the data-driven BT algorithms proposed in this paper are non-intrusive but they perform intrusive MOR on the interpolant of $G(s)$, for which a state-space realization can be conveniently obtained non-intrusively within the Loewner framework. Their non-intrusive nature does not stem from the fact that Hankel singular values are transfer function parameters and independent of a specific state-space realization, as argued in [11]. Instead, the balancing square-root algorithm remains intrinsically intrusive. Even in QuadBT, it operates on the intrusive state-space realization of the interpolant of $G(s)$, which interpolates $G(s)$ at the quadrature nodes.
@@
    The data-driven IRKA algorithms presented in this paper leverage the same principles as low-rank BT. They compute an interpolant of $G(s)$ by interpolating at several points to capture the majority of the dynamics of $G(s)$. This interpolant then serves as a surrogate for $G(s)$, allowing the algorithms to sample the interpolant as IRKA updates the interpolation points, rather than directly sampling $G(s)$. This approach enables the data-driven IRKA algorithms to bypass the need for new experiments to obtain additional samples of $G(s)$.
Having established that all interpolatory low-rank BT algorithms essentially reduce their respective data quadruplets, one might consider directly applying standard MOR algorithms like BT and IRKA to these quadruplets to obtain a compact ROM. However, these quadruplets are often not as well-behaved as desired. In many cases, when constructing an interpolant in the Loewner framework with a large number of interpolation points, the resulting interpolant is an unstable system with several poles in the right-half plane [7, 23]. As a result, standard MOR algorithms that require a stable original model cannot be directly applied to reduce the size of these quadruplets. Additionally, the Loewner matrix $W^{T}EV$ tends to become singular as the number of interpolation points increases [7, 23], rendering MOR algorithms that assume the non-singularity of the $E$-matrix unsuitable for directly reducing the order of Loewner interpolants. QuadBT and the algorithms proposed in this paper can be viewed as “compression” algorithms, designed to extract a compact ROM from these quadruplets. Moreover, these algorithms can also be seen as “distillation” algorithms, as they can extract ROMs with various properties from the same “raw” quadruplet by processing it differently. They effectively distill a compact, useful, and well-behaved ROM from the raw data quadruplets, which cannot be directly handled by standard MOR algorithms that assume the original model is well-behaved (like stable and minimal). This opens an interesting avenue for future research: developing similar compression strategies in intrusive settings to handle original systems that are similarly ill-behaved, much like these data quadruplets.
@@ 10 Numerical Examples
This section evaluates the performance of the proposed data-driven algorithms by comparing them to their intrusive counterparts. The effectiveness of these algorithms is illustrated through four numerical examples: the first two involve continuous-time systems, while the remaining two focus on discrete-time cases.
@@ 10.1 Experimental Setup
For quadrature-based algorithms, QuadBT [11] is first used to generate ROMs. The proposed quadrature-based IRKA algorithms then compress and distill the same quadruplets produced by QuadBT to target a local optimum of $||G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$. To ensure a fair comparison between DD-ADI-BT and frequency-domain QuadBT, the ADI shifts are chosen using (33) with $\zeta_{i,\sigma}=\zeta_{i,\mu}=10^{-4}$, so that both methods compress nearly identical quadruplets. Likewise, DD-ADI-BT and PORK-IRKA compress the same quadruplet, with $L_{p}$ and $L_{q}$ in DD-ADI-BT, and $Q_{s}^{-1}$ and $P_{s}^{-1}$ in PORK-IRKA, computed via (34)–(37).
For discrete-time systems, the procedure is similar: QuadBT [11] generates initial ROMs, and the proposed quadrature-based DT-IRKA algorithms compress and distill the same quadruplet to target a local optimum of $||G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$. To ensure a fair comparison between discrete-time QuadBT and DD-PORK-DTBT, interpolation points are selected using (96) with $\zeta_{i,\sigma}=\zeta_{i,\mu}=10^{-4}$, ensuring both methods act on nearly identical quadruplets. In this case, $L_{p}$ and $L_{q}$ in DD-PORK-DTBT, and $\bar{Q}_{s}^{-1}$ and $\bar{P}_{s}^{-1}$ in PORK-DTIRKA, are obtained from (97)–(100).
All IRKA and DT-IRKA-based algorithms are initialized arbitrarily and stopped after 50 iterations if convergence is not achieved. All experiments are conducted in MATLAB R2021b on a Windows 11 laptop with a 2GHz Intel Core i7 processor and 16GB of RAM.
@@ 10.2 CD Player
Figure 2 shows the largest 20 Hankel singular values approximated by QuadBT and DD-ADI-BT.
@@
As shown, DD-ADI-BT provides a close approximation to the Hankel singular values of the original system. Figure 3 presents the $\mathcal{H}_{\infty}$ norm of the relative error $\frac{||G(s)-G_{r}(s)||_{\mathcal{H}_{\infty}}}{||G(s)||_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1 through 20. The results indicate that DD-ADI-BT achieves accuracy comparable to that of intrusive BT and QuadBT.
Figure 3: Comparison of relative error $\frac{||G(s)-G_{r}(s)||_{\mathcal{H}_{\infty}}}{||G(s)||_{\mathcal{H}_{\infty}}}$
@@ 10.3 International Space Station
Figure 5 shows the 30 largest Hankel singular values approximated by QuadBT and DD-ADI-BT.
@@
As shown, DD-ADI-BT closely matches the Hankel singular values of the original system. Figure 6 presents the $\mathcal{H}_{\infty}$ norm of the relative error $\frac{||G(s)-G_{r}(s)||_{\mathcal{H}_{\infty}}}{||G(s)||_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1 through 30. The results demonstrate that DD-ADI-BT attains accuracy comparable to that of QuadBT.
Figure 6: Comparison of relative error $\frac{||G(s)-G_{r}(s)||_{\mathcal{H}_{\infty}}}{||G(s)||_{\mathcal{H}_{\infty}}}$
@@ 10.4 Low-pass Butterworth Filter
In this example, we consider the model from [11], which is a $40^{\text{th}}$-order low-pass Butterworth filter with a cutoff frequency of $0.6$ rad/sec. For frequency-domain QuadBT, 100 uniformly-spaced Gauss–Legendre quadrature nodes and weights are generated over the frequency range $-\pi$ to $\pi$ rad/sec. These are used to approximate both the controllability and observability Gramians. Transfer function samples at the quadrature nodes are obtained from the state-space realization of the Butterworth filter, generated using MATLAB’s ‘butter‘ command. For time-limited QuadBT, 100 impulse response samples are generated from the same state-space model. The corresponding quadruplets are then constructed and used to perform QuadBT. Next, 100 PORK shifts are computed as described in Subsection 10.1, followed by the evaluation of transfer function samples and construction of the associated quadruplet.
Figure 8 presents the 20 largest Hankel singular values approximated by QuadBT and DD-PORK-DTBT.
@@
As shown, DD-PORK-DTBT closely matches all 20 Hankel singular values. Figure 9 shows the $\mathcal{H}_{\infty}$ norm of the relative error $\frac{||G(z)-G_{r}(z)||_{\mathcal{H}_{\infty}}}{||G(z)||_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1 through 20. The results indicate that DD-PORK-DTBT provides accuracy comparable to that of QuadBT.
Figure 9: Comparison of relative error $\frac{||G(z)-G_{r}(z)||_{\mathcal{H}_{\infty}}}{||G(z)||_{\mathcal{H}_{\infty}}}$
@@ 10.5 Heat Transfer Equation
The Heat Transfer Equation model is a $200^{\text{th}}$-order system, taken from the benchmark collection in [44]. For frequency-domain QuadBT, 100 uniformly-spaced Gauss–Legendre quadrature nodes and weights are generated over the frequency range $-\pi$ to $\pi$ rad/sec. These are used to approximate both the controllability and observability Gramians. Transfer function samples at the quadrature nodes are obtained from the state-space realization provided in [44]. For time-limited QuadBT, 100 impulse response samples are generated from the same state-space model. The corresponding quadruplets are then constructed and used to perform QuadBT. Next, 100 PORK shifts are computed as described in Subsection 10.1, followed by the evaluation of transfer function samples and construction of the associated quadruplet.
Figure 11 presents the 30 largest Hankel singular values approximated by QuadBT and DD-PORK-DTBT.
@@
As shown, DD-PORK-DTBT closely most of the Hankel singular values. Figure 12 shows the $\mathcal{H}_{\infty}$ norm of the relative error $\frac{||G(z)-G_{r}(z)||_{\mathcal{H}_{\infty}}}{||G(z)||_{\mathcal{H}_{\infty}}}$ for ROMs of orders 1 through 30. The results indicate that DD-PORK-DTBT provides accuracy comparable to that of QuadBT.
Figure 12: Comparison of relative error $\frac{||G(z)-G_{r}(z)||_{\mathcal{H}_{\infty}}}{||G(z)||_{\mathcal{H}_{\infty}}}$
@@ 11 Conclusion
This paper presents data-driven, non-intrusive implementations of BT and IRKA for both continuous-time and discrete-time systems. The proposed methods utilize available frequency or time-domain data to compute ROMs. It has been observed that both QuadBT and the algorithms introduced in this paper effectively compress and distill their respective raw quadruplets, resulting in compact and practical ROMs. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
Appendix
A =[-3.2556 1.1671;-1.1671 -0.4906]; B =[-1.1810; -0.1793];
C =[1.1810 -0.1793]; n=size(A,1); E=eye(n);
a=[2.6141 1.1321]; % ADI Shifts
S = [2.6141 5.2282 3.4406 3.4406; 0 2.6141 3.4406 3.4406;
     0 0 1.1321 2.2642; 0 0 0 1.1321];
L = [-2.2865 -2.2865 -1.5047 -1.5047]; ns=size(S,1);
if rank(obsv(S,L)) == ns
    disp(’The pair (S,L) is observable’)
else
    disp(’Error: The pair (S,L) is not observable’)
end
w0=[1;2;3;4];
if rank(ctrb(S,w0)) == ns
    disp(’The pair (S,w(0)) is controllable’)
else
    disp(’Error: The pair (S,w(0)) is not controllable’)
end
sysw=ss(S,zeros(ns,1),L,0); % Signal Generator
tf=10; tspan=0:0.01:tf;
[u,t,w] = lsim(sysw,0.*tspan,tspan,w0); w=w.’; % Input Signal
sys=ss(A,B,C,0); x0=zeros(n,1);
[y,t,~] = lsim(sys,u,t,x0); % Output Data Generation
%Selected Input-Output Data Samples
o=n*ns+450; l=120; W=[]; Y=[];
for h=1:o
    ti=l+h;
    W=[W; w(:,ti).’]; Y=[Y; y(ti,:).’];
end
if rank(W) == ns
    disp(’Rank condition is satisfied’)
else
    disp(’Error: Move the data points or Increase the time window’)
end
At=-S’; Bt=L’; Ct=Y.’*W/(W.’*W); Et=eye(size(At)); % Interpolant
%Data Generation From Interpolant
[…]
%Loewner Quadruplet
Es=[-G_td_a1 -(Gt_a2-Gt_a1)/(a(2)-a(1));
    -(Gt_a1-Gt_a2)/(a(1)-a(2)) -G_td_a2];
As=[-Gt_a1-a(1)*G_td_a1 -(a(2)*Gt_a2-a(1)*Gt_a1)/(a(2)-a(1));
    -(a(1)*Gt_a1-a(2)*Gt_a2)/(a(1)-a(2)) -Gt_a2-a(2)*G_td_a2];
Bs=[Gt_a1; Gt_a2]; Cs=[Gt_a1 Gt_a2];
%DD-ADI-BT
Sb=diag([a(1) a(2)]); Lb=[1 1];  Sc=diag([a(1) a(2)]); Lc=[1;1];
Qs=lyap(-Sb’,Lb’*Lb); Lp=(chol(inv(Qs)))’;
Ps=lyap(-Sc,Lc*Lc’); Lq=(chol(inv(Ps)))’;
[Z_bsq,S_bsq,Y_bsq]=svd(full(Lq’*Es*Lp));
W=(Lq*Z_bsq)/(sqrtm(S_bsq)); V=(Lp*Y_bsq)/(sqrtm(S_bsq));
Ar=W’*As*V; Br=W’*Bs; Cr=Cs*V; sysr=ss(Ar,Br,Cr,0);
hsvd(sys)
hsvd(sysr)

### v4

@@ 8 Non-intrusive PORK-based DT-IRKA for Discrete Time Systems
It is now clear that DT-IRKA can be implemented using available transfer function samples $G(\alpha_{i})$ and $G(\beta_{i})$, eliminating the need for repeated estimations of $G(\sigma_{i})$ and $G^{\prime}(\sigma_{i})$ whenever DT-IRKA updates $\sigma_{i}$. The pseudo-code for the PORK-based DT-IRKA (PORK-DTIRKA) is outlined in Algorithm 8.
@@
2:while(relative change in $\lambda_{i}$ ¿ tol)
@@
6: Update the interpolation data: $(\sigma_{1},\cdots,\sigma_{r})=(\frac{1}{\lambda_{1}},\cdots,\frac{1}{\lambda_{r}})$; $b_{1}\cdots b_{r}]=B_{r}^{*}E_{r}^{-*}T_{r}^{-*}$; $c_{1}^{*}\cdots c_{r}^{*}]=C_{r}T_{r}$.
@@
8.1 Tracking the Error $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$
@@
[…]
Thus, with a delay of one iteration, the error $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}$ can be tracked by computing $G_{r}(z)||_{\mathcal{H}_{2}}^{2}$ in each iteration. Specifically, the variable component of the error in non-intrusive DT-IRKA can be monitored non-intrusively by tracking the following term:
[…]
However, it is important to note that the term $2\,\text{trace}\Big(C_{r}^{(i)}\big(C_{r}^{(i-1)}T_{r}^{(i-1)}\big)^{*}\Big)$ is an approximation and not exact. Its accuracy depends on the precision of the approximation of the integral (74) and (75) or the approximation of the infinite summation (84) and (85).
@@ 9 Compression and Distillation of Data Quadruplets
Throughout this paper, a consistent pattern has emerged in all the discussed non-intrusive algorithms: each algorithm constructs a Loewner quadruplet (in the frequency domain) or an impulse data quadruplet (in the time domain) and then reduces the respective data quadruplet, as illustrated in Figure 1.
@@
    Similar to numerical integration, interpolatory low-rank BT does not reduce $G(s)$ directly. Instead, an interpolant of $G(s)$ is first implicitly constructed (or explicitly constructed in non-intrusive settings). This interpolant is not particularly compact, as it is constructed to interpolate $G(s)$ at several interpolation points to capture the majority of the original system’s dynamics. Subsequently, this interpolant acts as a surrogate for $G(s)$. The ROMs produced by these low-rank BT algorithms are approximations of the interpolants of $G(s)$, rather than $G(s)$ itself. In this sense, low-rank BT could be termed “numerical BT” if we wish to adopt terminology analogous to numerical integration.
@@
    QuadBT and the non-intrusive BT algorithms proposed in this paper are non-intrusive but they perform intrusive MOR on the interpolant of $G(s)$, for which a state-space realization can be conveniently obtained non-intrusively within the Loewner framework.
@@
    The non-intrusive IRKA algorithms presented in this paper leverage the same principles as low-rank BT. They compute an interpolant of $G(s)$ by interpolating at several points to capture the majority of the dynamics of $G(s)$. This interpolant then serves as a surrogate for $G(s)$, allowing the algorithms to sample the interpolant as IRKA updates the interpolation points, rather than directly sampling $G(s)$. This approach enables the non-intrusive IRKA algorithms to bypass the need for new experiments to obtain additional samples of $G(s)$.
Having established that all interpolatory low-rank BT algorithms essentially reduce their respective data quadruplets, one might consider directly applying standard MOR algorithms like BT and IRKA to these quadruplets to obtain a compact ROM. However, these quadruplets are often not as well-behaved as desired. In many cases, when constructing an interpolant in the Loewner framework with a large number of interpolation points, the resulting interpolant is an unstable system with several poles in the right-half plane [7, 43]. As a result, standard MOR algorithms that require a stable original model cannot be directly applied to reduce the size of these quadruplets. Additionally, the Loewner matrix $W^{T}EV$ tends to become singular as the number of interpolation points increases [7, 43], rendering MOR algorithms that assume the non-singularity of the $E$-matrix unsuitable for directly reducing the order of Loewner interpolants. QuadBT and the algorithms proposed in this paper can be viewed as “compression” algorithms, designed to extract a compact ROM from these quadruplets. Moreover, these algorithms can also be seen as “distillation” algorithms, as they can extract ROMs with various properties from the same “raw” quadruplet by processing it differently. They effectively distill a compact, useful, and well-behaved ROM from the raw data quadruplets, which cannot be directly handled by standard MOR algorithms that assume the original model is well-behaved (like stable and minimal).
@@ 10 Numerical Examples
This section evaluates the performance of the proposed non-intrusive algorithms by comparing them to their intrusive counterparts. The effectiveness of these algorithms is illustrated through four numerical examples: the first two involve continuous-time systems, while the remaining two focus on discrete-time cases.
@@ 10.1 Experimental Setup
For quadrature-based algorithms, QuadBT [11] is first used to generate ROMs. The proposed quadrature-based IRKA algorithms then compress and distill the same quadruplets produced by QuadBT to target a local optimum of $G(s)-G_{r}(s)||_{\mathcal{H}_{2}}^{2}$. To ensure a fair comparison between NI-ADI-BT and frequency-domain QuadBT, the ADI shifts are chosen using (33) with $\zeta_{i,\sigma}=\zeta_{i,\mu}=10^{-4}$, so that both methods compress nearly identical quadruplets. Likewise, NI-ADI-BT and PORK-IRKA compress the same quadruplet, with $L_{p}$ and $L_{q}$ in NI-ADI-BT, and $Q_{s}^{-1}$ and $P_{s}^{-1}$ in PORK-IRKA, computed via (34)–(37).
For discrete-time systems, the procedure is similar: QuadBT [11] generates initial ROMs, and the proposed quadrature-based DT-IRKA algorithms compress and distill the same quadruplet to target a local optimum of $G(z)-G_{r}(z)||_{\mathcal{H}_{2}}^{2}$. To ensure a fair comparison between discrete-time QuadBT and NI-PORK-DTBT, interpolation points are selected using (97) with $\zeta_{i,\sigma}=\zeta_{i,\mu}=10^{-4}$, ensuring both methods act on nearly identical quadruplets. In this case, $L_{p}$ and $L_{q}$ in NI-PORK-DTBT, and $\bar{Q}_{s}^{-1}$ and $\bar{P}_{s}^{-1}$ in PORK-DTIRKA, are obtained from (98)–(101).
All IRKA and DT-IRKA-based algorithms are initialized arbitrarily and stopped after 50 iterations if convergence is not achieved. All experiments are conducted in MATLAB R2021b on a Windows 11 laptop with a 2GHz Intel Core i7 processor and 16GB of RAM. The MATLAB codes for reproducing the results in this section are provided in [42].
@@ 10.2 CD Player
Figure 2 shows the largest 20 Hankel singular values approximated by QuadBT and NI-ADI-BT.
@@
[…]
@@ 10.3 International Space Station
Figure 5 shows the 30 largest Hankel singular values approximated by QuadBT and NI-ADI-BT.
@@
[…]
@@ 10.4 Low-pass Butterworth Filter
In this example, we consider the model from [11], which is a $40^{\text{th}}$-order low-pass Butterworth filter with a cutoff frequency of $0.6$ rad/sec. For frequency-domain QuadBT, 100 uniformly-spaced Gauss–Legendre quadrature nodes and weights are generated over the frequency range $\pi$ to $\pi$ rad/sec. These are used to approximate both the controllability and observability Gramians. Transfer function samples at the quadrature nodes are obtained from the state-space realization of the Butterworth filter, generated using MATLAB’s ‘butter‘ command. For time-limited QuadBT, 100 impulse response samples are generated from the same state-space model. The corresponding quadruplets are then constructed and used to perform QuadBT. Next, 100 PORK shifts are computed as described in Subsection 10.1, followed by the evaluation of transfer function samples and construction of the associated quadruplet.
Figure 8 presents the 20 largest Hankel singular values approximated by QuadBT and NI-PORK-DTBT.
@@
[…]
@@ 10.5 Heat Transfer Equation
The Heat Transfer Equation model is a $200^{\text{th}}$-order system, taken from the benchmark collection in [44]. For frequency-domain QuadBT, 100 uniformly-spaced Gauss–Legendre quadrature nodes and weights are generated over the frequency range $\pi$ to $\pi$ rad/sec. These are used to approximate both the controllability and observability Gramians. Transfer function samples at the quadrature nodes are obtained from the state-space realization provided in [44]. For time-limited QuadBT, 100 impulse response samples are generated from the same state-space model. The corresponding quadruplets are then constructed and used to perform QuadBT. Next, 100 PORK shifts are computed as described in Subsection 10.1, followed by the evaluation of transfer function samples and construction of the associated quadruplet.
Figure 11 presents the 30 largest Hankel singular values approximated by QuadBT and NI-PORK-DTBT.
@@
[…]
@@ 11 Conclusion
This paper presents non-intrusive implementations of BT and IRKA for both continuous-time and discrete-time systems. The proposed methods utilize available frequency or time-domain data to compute ROMs. It has been observed that both QuadBT and the algorithms introduced in this paper effectively compress and distill their respective raw quadruplets, resulting in compact and practical ROMs. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
Acknowledgements and Funding
We are deeply grateful to Ion Victor Gosea at the Max Planck Institute for Dynamics of Complex Technical Systems in Magdeburg, Germany, for his patient responses to our numerous questions about the Loewner framework and for his valuable feedback that helped improve our manuscript. This work was supported by the National Natural Science Foundation of China under Grants No. 62350410484, 62273059, 62376146, and 62176001, and in part by the Shanxi Province Central Guidance for Local Science and Technology Development Special Project (Grant No. YDZJSX20231D003).